OpenStudy (anonymous):

find the 6 trig functions of cos theta=.354 . I have no idea what to do with the decimal number. Where on the triangle does it go? I'm not even sure where to start.

OpenStudy (anonymous):

SO, do I just turn .354 into a fraction: 354/1000 and reduce it down to a manageable number?

OpenStudy (anonymous):

So if it is like 2.5 would be 2 1/2 or 25/10?

OpenStudy (anonymous):

\[\text{Sin}\left[\text{ArcCos}\left[\frac{177}{500}\right]\right]=\frac{\sqrt{218671}}{500}=0.935245 \]

OpenStudy (anonymous):

\[\text{ArcCos}\left[\frac{177}{500}\right]=1.20895=69.2677{}^{\circ} \] Using the degree result will allow you to find all six trig function values.

OpenStudy (anonymous):

By the way,\[\frac{354}{1000}=\frac{177}{500} \]